OpenStudy (anonymous):

In exercises 12 and 13, the midpoints of the sides of triangle HJK are L(4,3), M(8,3), and N(6,1). A) Find the coordinates of the vertices of the triangle B) Show that each midsegment is parallel to a side of the triangle.

OpenStudy (anonymous):

Can anybody please help? I constructed the given points, but I am not sure how to find the coordinates of the vertices of the triangle

OpenStudy (anonymous):

is answer (4,1), (4,5), (8,1)

OpenStudy (anonymous):

@bhusal can you please explain? :( I don't understand

OpenStudy (anonymous):

@RH any problem

OpenStudy (anonymous):

@bhusal I don't understand what you did in the beginning :( sorry

OpenStudy (anonymous):

\[i supposed the coordinate of vertices as (x1,y1), (x2,y2), (x3,y3) and i used midpoint formulai.e \mid-points (x,y)=(x1+x2\div2 , Y1+y2\div2)\]

OpenStudy (anonymous):

i supposed the coordinate of vertices as (x1,y1), (x2,y2), (x3,y3) and i used midpoint formula i.e mid-points (x,y)=(x1+x2/2 , Y1+y2/2)

OpenStudy (anonymous):

did you get it

OpenStudy (anonymous):

no!! :( I am sorry leave it I can't understand at all what you did ! :( I am sorry to take your time

OpenStudy (anonymous):

dont feel sorry where did you stuck is i shud help

OpenStudy (anonymous):

do you have knowlwdge on midpoints of line segment

OpenStudy (anonymous):

Yes I know what midpoints are. I don't understand how you find the points. so the given points are the midpoints of the triangle. In my book it says that I should find the slopes to draw the triangle. that is what I don't understand. How to find the points.

OpenStudy (anonymous):

hey @RH did you solved

OpenStudy (anonymous):

I think I understand. Are the points for HJ (8.6) ?

OpenStudy (anonymous):

what do you want to say by points for HJ

OpenStudy (anonymous):

is point H(8,6) ?

OpenStudy (anonymous):

i got (4,1) its all about solving i may have done mistake you have understood the theme of question thats enough

OpenStudy (anonymous):

I think I did. So you always find the slope first and then solve right? or..

OpenStudy (anonymous):

know i find points first slope= y2-y1/x2-x1 for that i need x1,y1 and x2,y2